OpenStudy (anonymous):

A scientist counts 35 bacteria present in a culture and finds that the number of bacteria triples each hour. The function y = 35 • 3x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (anonymous):

you have to devide 3 from both sides

OpenStudy (anonymous):

y = 35 • 3^x 550 = 35 • 3^x Approx. 15.7143 = 3^x

OpenStudy (anonymous):

@Najou

OpenStudy (anonymous):

i dont know this i'm in 8th grade sorry

OpenStudy (anonymous):

no i said devide 35 by three

OpenStudy (anonymous):

that;s ur 1st stp